Priyanka's age is thrice her age three years hence minus her age three years ago. what is her present age.
Answer by fractalier(6550)

Call her age now, p. Three years hence is p+3. Three years ago is p-3.
p = 3(p+3) - (p-3)
p = 3p + 9 - p + 3
p = 2p + 12
-p = 12
p = -12
Obviously there is something wrong with the problem, unless it is
p = 3[(p+3) - (p-3)]
p = 3(6) = 18
If so, the problem isn't worded precisely.